# You Only Learn One Query: Learning Unified Human Query for Single-Stage Multi-Person Multi-Task Human-Centric Perception

### v2

@@ Abstract
Human-centric perception (e.g. pedetrian detection, segmentation, pose estimation, and attribute analysis) is a long-standing problem for computer vision. This paper introduces a unified and versatile framework (HQNet) for single-stage multi-person multi-task human-centric perception (HCP). Our approach centers on learning a unified human query representation, denoted as Human Query, which captures intricate instance-level features for individual persons and disentangles complex multi-person scenarios. Although different HCP tasks have been well-studied individually, single-stage multi-task learning of HCP tasks has not been fully exploited in the literature due to the absence of a comprehensive benchmark dataset. To address this gap, we propose COCO-UniHuman benchmark dataset to enable model development and comprehensive evaluation. Experimental results demonstrate the proposed method’s state-of-the-art performance among multi-task HCP models and its competitive performance compared to task-specific HCP models. Moreover, our experiments underscore Human Query’s adaptability to new HCP tasks, thus demonstrating its robust generalization capability. Codes and data will be publicly accessible.
@@ 1 Introduction
Human-centric visual perception (e.g. pedestrian detection, 2D keypoint estimation, 3D mesh recovery, human segmentation and attribute recognition) have attracted increasing research attention owing to their widespread industrial applications such as sports analysis, virtual reality, and augmented reality.
@@
2 Related Works
2.1 Human-Centric Perception Tasks and Datasets
Human-centric perception (HCP) tasks including keypoint estimation, segmentation, and attribute recognition, have been extensively studied in computer vision. Approaches to multi-person HCP can be categorized into top-down, bottom-up, and single-stage methods. Top-down methods follow a detect-then-analyze approach. They first localize human instances, and then perform single person analysis. Top-down approaches can be divided into two types: those using separate pre-trained detectors and task-specific perception models [63, 30, 84, 42, 67, 74], and those jointly learning detection and perception modules [20]. Bottom-up methods learn instance-agnostic keypoints/masks and cluster them using integer linear programming [23, 32], heuristic greedy parsing [6, 59], embedding clustering [54, 33], or learnable clustering [26]. Single-stage methods directly predict keypoints or masks for each individual, with different representations for 2D keypoint estimation (coordinate-based [71, 56, 76, 81], heatmap-based [73, 66], or hybrid [94, 53, 18]), 3D mesh recovery [68, 41] and segmentation (contour-based [79] or mask-based [5]). While existing approaches focus on individual HCP tasks, we aim to unify HCP by learning a single model that handles multiple tasks simultaneously, enabling a comprehensive understanding of humans.
Table 1: Overview of representative HCP datasets. “# Img”, “# Inst”, and “# ID” mean the number of total images, instances and identities respectively. “Crop” indicates whether the images are cropped for “face” or “body”. “*” means head box annotation. “group:n” means age classification with n groups, “real” means real age estimation, and “appa” means apparent age estimation.
@@ 2.1 Human-Centric Perception Tasks and Datasets
As shown in Table 1, there are many task-specific datasets separately annotated for different HCP tasks, including pedestrian detection [15, 88, 64], multi-human keypoint estimation [4, 3], multi-human segmentation [19, 37], and human attribute recognition [92, 49]. Datasets for multiple HCP tasks also exist. Widely used COCO [45, 27] offers thorough annotations: body box, keypoints, and segmentation mask. Our COCO-UniHuman dataset further extends COCO featuring extensive gender and age annotations.
@@ 2.2 Unified Methods for HCP
General network architecture for different HCP tasks. Some works design general network backbones, including CNN-based [74] and Transformer-based backbones [83]. Others unify HCP tasks with novel perception heads, such as UniHead [40]. Unlike these methods, which employ separate task-specific models, we consolidate diverse HCP tasks within a single network. Pre-training on HCP tasks. There are also works [22, 9, 69] on pre-training on diverse human-centric tasks with large-scale data. More recently, UniHCP [11] presents a unified vision transformer model to perform multitask pre-training at scale. It employs task-specific queries for attending to relevant features, but tackles one task at a time. Unlike ours, our approach simultaneously solves multiple HCP tasks in a single forward pass. Our approach contrasts with these pre-training based methods by avoiding pre-training, minimizing fine-tuning, and circumventing resource-intensive multi-dataset training. Unlike them, we handle multiple HCP tasks concurrently in a single-stage, multi-task manner, diverging from their single-person focus. Co-learning on HCP tasks. Many works have investigated the correlations between pairs of HCP tasks [86, 70, 46, 55, 57]. We propose a single-stage model that learns a general unified representation to handle all representative human-centric perception tasks simultaneously.
@@ 3 COCO-UniHuman Dataset
COCO-UniHuman is the first large-scale dataset, which provides annotations for all four representative HCP tasks in multi-person scenarios. Building upon MSCOCO’2017 [45, 27, 28] dataset, we have enriched the annotations by including gender, age, 3D body mesh information for each individual.
@@ 3.1 Data Annotation
To ensure accurate annotations, we employ trained annotators to manually label the gender and apparent age for each human instance in the dataset. We discard images full of non-human objects, and exclude all Small category persons that are hardly attribute-recognizable. Gender annotation. For each valid human instance, we adopt a body-based annotation approach. Using the provided human bounding boxes, we crop the body images and request annotators to label the gender. To maintain data quality, we conduct quality inspections and manual corrections throughout the labeling process. Age annotation. To enhance the quality of annotation, we employ a two-stage strategy based on body-based annotation. Similar to gender annotation, age annotation is also performed on cropped body images. We implement a coarse-to-fine two-stage annotation strategy, considering age group annotation to be comparatively easier than apparent age annotation [2]. In the first stage, age groups are annotated. Following [39], we divide the age ranges into six groups, i.e. “baby”, “kid”, “teen”, “young”, “middle aged”, and “elderly”. For each cropped person image, we request a group of 10 annotators to independently and repeatedly label the age groups (6-category classification task). We take the mode of the 10 votes as the ground-truth age group. In the second stage, the apparent age is annotated. Given the age group as a prior, a group of 10 annotators independently and repeatedly annotate the apparent age. Consequently, we obtain 10 votes for each human instance. We remove the outliers and take the average as the final ground-truth apparent age. As a summary, the dataset contains over 1M apparent votes. Experiments validate the effectiveness of the body-based annotation strategy (Sec. A2.1) and the two-stage annotation strategy (Sec. A2.2). Mesh Annotation. We follow [28] to apply Exemplar Fine-Tuning (EFT) method to generate 3D pseudo-ground truth SMPL parameters. To ensure data quality, we only retain instances with at least 12 keypoint annotations where all limbs are visible and filter out low-quality data manually.
@@ 4.1 Overview
This study endeavors to develop a single-stage framework that supports a wide range of human-centric perception (HCP) tasks. The key is to learn a comprehensive human representation, which can be universally employed across various HCP tasks. To achieve this, we employ a query-based methodology and investigate the feasibility of representing each human instance as a single shared query. Our objective is to build a framework that is characterized by simplicity, flexibility, and scalability. Unlike previous task-specific HCP models that may incorporate specialized designs tailored to specific tasks (e.g. “mask-enhanced anchor box initialization” in Mask DINO [36]), our approach aims to handle various human-centric analysis tasks in a unified manner. To maximize knowledge sharing among various HCP tasks, we attempt to share most weights across different HCP tasks.
@@ 4.2 Task-Shared Transformer Decoder
HumanQuery-Instance Matching. To ensure consistent and unique predictions for each ground-truth instance across all HCP tasks, including classification (Cls.), detection (Det.), pose (Pose.), and segmentation (Seg.), we employ HumanQuery-Instance (HQ-Ins) Matching. $\lambda_{cls}L_{cls}+\lambda_{det}L_{det}+\lambda_{seg}L_{seg}+\lambda_{pose}L_{pose},$ where $\lambda$ are the corresponding loss weights. Further details can be found in Sec. A4.1.
@@ 4.3 Task-Specific Heads
To minimize the overhead of incorporating new tasks, we employ lightweight task-specific heads comprising simple linear layers. The details of our four task-specific heads are as follows: Human detection head. A 3-layer multi-layer perceptron (MLP) with a hidden dimension of $d$ is utilized to predict the normalized center x-y coordinates, height, and width of the bounding box w.r.t. the input image. Additionally, a linear projection layer (FC) is employed to predict the class label (human or non-human). 2D keypoint estimation head. Following the coordinate prediction paradigm, the learned Human Query is fed into a pose regression head (MLP) to regress the relative pose offsets w.r.t. the shared reference points of the detection head. A confidence prediction head (FC) is used to predict confidence score of having visible keypoints. Following PETR [65], joint decoder layers are employed to refine body poses by leveraging structured relations between body keypoints. An auxiliary heatmap branch is used to aid training and discarded during testing. Human instance segmentation head. A 3-layer MLP is used to process the instance-aware pixel embedding map and output a one-channel mask, which is then upsampled to match the original input image size. Human attribute head. The gender estimation head and the age estimation head operate in parallel. Both heads consist of two-layer MLPs. Gender estimation involves binary classification, while age estimation is formulated as an 85-class ([1, 85]) classification with softmax expected value [62] estimation. 3D mesh recovery head. Two 3-layer MLPs with the same hidden dimension are used to predict the pose and shape parameters respectively. These parameters are then fed into a SMPL body model to generate the 3D body meshes.
@@ 5 Experiments
$\diamondsuit$ ${}^{\dagger}$ [78] R-50 70.4 67.1 77.2
$\diamondsuit$ ${}^{\dagger}$ [50] Swin-L 74.3 70.6 81.2
$\diamondsuit$ ${}^{\dagger}$ [67] HRNet-32 74.4 70.8 81.0
$\diamondsuit$ ${}^{\dagger}$ [38] R-50 68.2 63.2 76.2
$\heartsuit$ ${}^{\dagger}$ [10] HRNet-w32 67.1 61.5 76.1
$\heartsuit$ ${}^{\dagger}$ [18] HRNet-w32 68.0 62.1 77.7
$\heartsuit$ ${}^{\dagger}$ [52] HRNet-w32 68.9 63.0 77.4
$\bigstar$ ${}^{\dagger}$ [73] R-50-FPN 52.0 48.6 58.0
$\bigstar$ ${}^{\dagger}$ [73] HRNet-w32 69.8 64.0 78.9
@@
Model Backbone Pose (Mesh.)
@@
ROMP [68] R-50 119.52 72.27
HQNet w/o GaSM R-50 87.00 54.92
HQNet R-50 84.74 50.80
HQNet ViT-L 76.31 48.26
5.1 Dataset and Evaluation Metric
COCO-UniHuman Dataset. Our model training exclusively employs COCO-UniHuman train data (in addition to ImageNet pre-training). We follow DINO [85] for augmentation and adopt the 100-epoch training schedule. Model evaluation takes place on COCO-UniHuman val set (2693 images). Due to limitations of the COCO test-dev evaluation server, which lacks support for “Person" category evaluation and attribute recognition, we mainly report results on the val set. The evaluation is based on the standard COCO metrics including Average Precision (AP), $\text{AP}^{M}$ for medium-sized persons and $\text{AP}^{L}$ for large-sized persons. Following [89, 90], we exclude the Small category persons during evaluation due to the lack of annotations in COCO. For attribute recognition, we also use AP with Age-10 metric for evaluation, where the age estimation is considered correct if the prediction error is no larger than 10. For human mesh recovery, we evaluate pose accuracy using MPJPE (Mean Per Joint Position Error) w.r.t. root relative poses and PA-MPJPE (Procrustes-Aligned MPJPE), which is MPJPE calculated after rigid alignment of predicted pose with the ground truth.
@@ 5.2 Results on COCO-UniHuman Dataset
We compare our method to task-specific and multi-task HCP models on the COCO-UniHuman dataset in Table 2. Our models outperform multi-task HCP models and achieves very competitive results against task-specific HCP models. Details about the baselines can be found in Sec. A5. “D”, “S”, “P”, “C” mean model training with Detection (Det.), Segmentation (Seg.), Pose and Classification (Cls.) task respectively.
Comparison with task-specific HCP models. For human detection, we compare two baseline approaches, i.e. Faster-RCNN [60] and DINO [85]. For human instance segmentation, we contrast HQNet with state-of-the-art general and human-specific instance segmentation methods, including Mask R-CNN [20], PolarMask [79], MEInst [87], YOLACT [5], and CondInst [72]. For human pose estimation, we compare with several representative top-down methods (SBL [78], HRNet [67], Swin [50] and PRTR [38]), bottom-up approaches (HrHRNet [10], DEKR [18], and SWAHR [52]) and single-stage approaches (FCPose [53], InsPose [66], PETR [65] and CID [73]). For gender and age estimation, we establish baselines using StrongBL [25] and Mask R-CNN [20]. Our approach achieves very competitive performance compared to other task-specific HCP models when using the R-50 backbone. Moreover, with stronger backbones such as Swin-L and ViT-L, we establish new state-of-the-art results.
@@
$\heartsuit$ ${}^{\dagger}$ [54] Hourglass 32.1 29.5
$\heartsuit$ ${}^{\dagger}$ [26] Hourglass 35.6 34.8
$\heartsuit$ ${}^{\dagger}$ [18] HRNet-w32 37.9 36.5
$\heartsuit$ ${}^{\dagger}$ [10] HRNet-w32 40.0 39.4
@@
$\bigstar$ ${}^{\dagger}$ [81] HRNet-w32 39.0 38.1
$\bigstar$ ${}^{\dagger}$ [73] R-50-FPN 29.2 28.3
$\bigstar$ ${}^{\dagger}$ [73] HRNet-w32 44.9 44.0

### v3

@@ Abstract
Human-centric perception (e.g. detection, segmentation, pose estimation, and attribute analysis) is a long-standing problem for computer vision. This paper introduces a unified and versatile framework (HQNet) for single-stage multi-person multi-task human-centric perception (HCP). Our approach centers on learning a unified human query representation, denoted as Human Query, which captures intricate instance-level features for individual persons and disentangles complex multi-person scenarios. Although different HCP tasks have been well-studied individually, single-stage multi-task learning of HCP tasks has not been fully exploited in the literature due to the absence of a comprehensive benchmark dataset. To address this gap, we propose COCO-UniHuman benchmark to enable model development and comprehensive evaluation. Experimental results demonstrate the proposed method’s state-of-the-art performance among multi-task HCP models and its competitive performance compared to task-specific HCP models. Moreover, our experiments underscore Human Query’s adaptability to new HCP tasks, thus demonstrating its robust generalization capability. Codes and data are available at https://github.com/lishuhuai527/COCO-UniHuman.
@@ 1 Introduction
Human-centric perception (e.g. pedestrian detection, 2D keypoint estimation, 3D mesh recovery, human segmentation and attribute recognition) have attracted increasing research attention owing to their widespread industrial applications such as sports analysis, virtual reality, and augmented reality.
@@
Table 1: Overview of representative HCP datasets. “#Img”, “#Inst”, and “#ID” mean the number of total images, instances and identities respectively. “Crop” indicates whether the images are cropped for “face” or “body”. * means head box annotation. “group:n” means age classification with n groups, “real” means real age estimation, and “appa” means apparent age estimation.
@@
2 Related Works
2.1 Human-Centric Perception Tasks and Datasets
Approaches to multi-person human-centric perception (HCP) can be categorized into top-down, bottom-up, and single-stage methods. Top-down methods follow a detect-then-analyze approach. They first localize human instances, and then perform single person analysis. Top-down approaches can be divided into two types: those using separate pre-trained detectors and task-specific perception models [34, 90, 46, 71, 78, 27, 84], and those jointly learning detection and perception modules [21, 3]. Bottom-up methods learn instance-agnostic keypoints/masks and cluster them using integer linear programming [24, 36, 28], heuristic greedy parsing [7, 64], embedding clustering [58, 37], or learnable clustering [29]. Single-stage methods directly predict keypoints or masks for each individual, with different representations for 2D keypoint estimation (coordinate-based [75, 60, 80, 87], heatmap-based [77, 70], or hybrid [102, 57, 19]), 3D mesh recovery [72, 45] and segmentation (contour-based [83] or mask-based [6]). While existing approaches focus on individual HCP tasks, we aim to unify HCP by learning a single model that handles multiple tasks simultaneously, enabling a comprehensive understanding of humans. As shown in Table 1, there are task-specific datasets separately annotated for different HCP tasks, including pedestrian detection [97, 16, 94, 68], keypoint estimation [5, 4], segmentation [20, 41], and attribute recognition [99, 53]. Datasets for multiple HCP tasks also exist. COCO [49] offers thorough annotations: body box, keypoints, and segmentation mask. COCO-WholeBody [30, 85] provides dense annotations of face/hand boxes and 133 whole-body keypoints. Our COCO-UniHuman dataset further extends COCO-WholeBody featuring extensive gender, age and mesh annotations.
@@ 2.2 Unified Methods for HCP
General network architecture for different HCP tasks. Some works design general network backbones, including CNN-based [78] and Transformer-based backbones [89]. Others unify HCP tasks with novel perception heads, such as UniHead [44] and UniFS [31]. Unlike these methods, which employ separate task-specific models, we consolidate diverse HCP tasks within a single network. Pre-training on HCP tasks. There are also works [23, 10, 73] on pre-training on diverse human-centric tasks with large-scale data. More recently, UniHCP [12] presents a unified vision transformer model to perform multitask pre-training at scale. It employs task-specific queries for attending to relevant features, but tackles one task at a time. Unlike ours, our approach simultaneously solves multiple HCP tasks in a single forward pass. Our approach contrasts with these pre-training based methods by avoiding pre-training, minimizing fine-tuning, and circumventing resource-intensive multi-dataset training. Unlike them, we handle multiple HCP tasks concurrently in a single-stage, multi-task manner, diverging from their single-person focus. Co-learning on HCP tasks. Many works have investigated the correlations between pairs of HCP tasks [92, 74, 50, 59, 61]. We propose a single-stage model that learns a general unified representation to handle all representative human-centric perception tasks simultaneously.
@@ 3 COCO-UniHuman Dataset
COCO-UniHuman v1 dataset is the first large-scale dataset, which provides annotations for all four representative HCP tasks in multi-person scenarios. Building upon COCO [49, 32] dataset, we have enriched the annotations by including gender, age, 3D body mesh information for each individual.
@@ 3.1 Data Annotation
To ensure accurate annotations, we employ trained annotators to manually label the gender and apparent age for each human instance in the dataset. We discard images full of non-human objects, and exclude all Small category persons that are hardly attribute-recognizable. Gender annotation. For each valid human instance, we adopt a body-based annotation approach. Using the provided human bounding boxes, we crop the body images and request annotators to label the gender. To maintain data quality, we conduct quality inspections and manual corrections throughout the labeling process. Age annotation. To enhance the quality of annotation, we employ a two-stage strategy based on body-based annotation. Similar to gender annotation, age annotation is also performed on cropped body images. We implement a coarse-to-fine two-stage annotation strategy, considering age group annotation to be comparatively easier than apparent age annotation [2]. In the first stage, age groups are annotated. Following [43], we divide the age ranges into six groups, i.e. “baby”, “kid”, “teen”, “young”, “middle aged”, and “elderly”. For each cropped person image, we request a group of 10 annotators to independently and repeatedly label the age groups (6-category classification task). We take the mode of the 10 votes as the ground-truth age group. In the second stage, the apparent age is annotated. Given the age group as a prior, a group of 10 annotators independently and repeatedly annotate the apparent age. Consequently, we obtain 10 votes for each human instance. We remove the outliers and take the average as the final ground-truth apparent age. As a summary, the dataset contains over 1M apparent votes. Experiments validate the effectiveness of the body-based annotation strategy and the two-stage annotation strategy (see Supplementary). Mesh Annotation. We follow [32] to apply Exemplar Fine-Tuning (EFT) method with the gender neutral model to generate 3D pseudo-ground truth SMPL parameters. To ensure data quality, we only retain instances with at least 12 keypoint annotations where all limbs are visible and filter out low-quality data manually.
@@ 4.1 Overview
This study endeavors to develop a single-stage framework that supports a wide range of human-centric perception (HCP) tasks. The key is to learn a comprehensive human representation, which can be universally employed across various HCP tasks. To achieve this, we employ a query-based methodology and investigate the feasibility of representing each human instance as a single shared query. Unlike previous task-specific HCP models that may incorporate specialized designs tailored to specific tasks (e.g. “mask-enhanced anchor box initialization” in Mask DINO [40]), our approach aims to handle various human-centric analysis tasks in a unified manner. To maximize knowledge sharing among various HCP tasks, we attempt to share most weights across different HCP tasks.
@@ 4.2 Task-Shared Transformer Decoder
HumanQuery-Instance Matching. To ensure consistent and unique predictions for each ground-truth instance across all HCP tasks, i.e. classification (Cls.), detection (Det.), pose (Pose.), and segmentation (Seg.), we employ HumanQuery-Instance (HQ-Ins) Matching. $\lambda_{cls}L_{cls}+\lambda_{det}L_{det}+\lambda_{seg}L_{seg}+\lambda_{pose}L_{pose},$ where $\lambda$ are loss weights. Details can be found in Supplementary.
@@ 4.3 Task-Specific Heads
To minimize the overhead of incorporating new tasks, we employ lightweight task-specific heads. Human detection head. A 3-layer multi-layer perceptron (MLP) with a hidden dimension of $d$ is utilized to predict the normalized center x-y coordinates, height, and width of the bounding box w.r.t. the input image. Additionally, a linear projection layer (FC) is employed to predict the class label (human or non-human). 2D keypoint estimation head. Following the coordinate prediction paradigm, the learned Human Query is fed into a pose regression head (MLP) to regress the relative pose offsets w.r.t. the shared reference points of the detection head. A confidence prediction head (FC) is used to predict confidence score of having visible keypoints. Following PETR [69], joint decoder layers are employed to refine body poses by leveraging structured relations between body keypoints. An auxiliary heatmap branch is used to aid training and discarded during testing. Human instance segmentation head. A 3-layer MLP is used to process the instance-aware pixel embedding map and output a one-channel mask, which is then upsampled to match the original input image size. Human attribute head. The gender estimation head and the age estimation head operate in parallel. Both heads consist of two-layer MLPs. Gender estimation involves binary classification, while age estimation is formulated as an 85-class ([1, 85]) classification with softmax expected value [67] estimation. 3D mesh recovery head. Two 3-layer MLPs with the same hidden dimension are used to predict the pose and shape parameters respectively. These parameters are then fed into a SMPL body model to generate the 3D body meshes.
@@ 5 Experiments
$\diamondsuit$ SBL [82] R-50 70.4 67.1 77.2
$\diamondsuit$ Swin [54] Swin-L 74.3 70.6 81.2
$\diamondsuit$ HRNet [71] HRNet-32 74.4 70.8 81.0
$\diamondsuit$ ViTPose [86] ViT-L 78.2 74.5 85.4
$\diamondsuit$ PRTR [42] R-50 68.2 63.2 76.2
$\heartsuit$ HrHRNet [11] HRNet-w32 67.1 61.5 76.1
$\heartsuit$ DEKR [19] HRNet-w32 68.0 62.1 77.7
$\heartsuit$ SWAHR [56] HRNet-w32 68.9 63.0 77.4
$\bigstar$ CID [77] R-50-FPN 52.0 48.6 58.0
$\bigstar$ CID [77] HRNet-w32 69.8 64.0 78.9
@@
5.1 Dataset and Evaluation Metric
COCO-UniHuman Dataset. Our model training exclusively employs COCO-UniHuman train data (in addition to ImageNet pre-training). We follow DINO [91] for augmentation and adopt the 100-epoch training schedule. Model evaluation takes place on COCO-UniHuman val set (2693 images). Due to limitations of the COCO test-dev evaluation server, which lacks support for “Person” category evaluation and attribute recognition, we mainly report results on the val set. The evaluation is based on the standard COCO metrics including Average Precision (AP), $\text{AP}^{M}$ for medium-sized persons and $\text{AP}^{L}$ for large-sized persons. Following [95, 96], we exclude the Small category persons during evaluation due to the lack of annotations in COCO. For attribute recognition, we also use AP with Age-10 metric for evaluation, where the age estimation is considered correct if the prediction error is no larger than 10. For human mesh recovery, we evaluate pose accuracy using MPJPE (Mean Per Joint Position Error) w.r.t. root relative poses and PA-MPJPE (Procrustes-Aligned MPJPE), which is MPJPE calculated after rigid alignment of predicted pose with the ground truth.
@@ 5.1 Dataset and Evaluation Metric
Model Backbone Bbox Pose (Mesh.)
@@
$\diamondsuit$ HMR [34] R-50 GT 109.62 72.03
$\diamondsuit$ HMR+ [63] R-50 GT 78.06 50.36
$\bigstar$ ROMP [72] R-50 - 119.52 72.27
$\bigstar$ HQNet w/o GaSM R-50 - 87.00 54.92
$\bigstar$ HQNet R-50 - 84.74 50.80
$\bigstar$ HQNet ViT-L - 76.31 48.26
@@ 5.2 Results on COCO-UniHuman Dataset
We compare our method to task-specific and multi-task HCP models on the COCO-UniHuman dataset in Table 2 and Table 3. Our models outperform multi-task HCP models and achieves very competitive results against task-specific HCP models. Details about the baselines can be found in Supplementary. “D”, “S”, “P”, “C” mean model training with Detection (Det.), Segmentation (Seg.), Pose and Classification (Cls.) task respectively.
Comparison with task-specific HCP models. For human detection, we compare three baseline approaches, i.e. Faster-RCNN [65], IterDETR [100] and DINO [91]. For human instance segmentation, we contrast HQNet with state-of-the-art general and human-specific instance segmentation methods, including Mask R-CNN [21], PolarMask [83], MEInst [93], YOLACT [6], and CondInst [76]. For human pose estimation, we compare with several representative top-down methods (SBL [82], HRNet [71], Swin [54], ViTPose [86] and PRTR [42]), bottom-up approaches (HrHRNet [11], DEKR [19], and SWAHR [56]) and single-stage approaches (FCPose [57], InsPose [70], PETR [69] and CID [77]). For gender and age estimation, we establish baselines using StrongBL [26] and Mask R-CNN [21]. For mesh, we compare with HMR [34], HMR+ [63] and ROMP [72]. Our approach achieves very competitive performance compared to other task-specific HCP models when using the R-50 backbone. Moreover, with stronger backbones such as Swin-L and ViT-L, we achieve SOTA among single-stage approaches.
@@
$\heartsuit$ AE [58] Hourglass 32.1 29.5
$\heartsuit$ HGG [29] Hourglass 35.6 34.8
$\heartsuit$ DEKR [19] HRNet-w32 37.9 36.5
$\heartsuit$ HrHRNet [11] HRNet-w32 40.0 39.4
@@
$\bigstar$ LOGO-CAP [87] HRNet-w32 39.0 38.1
$\bigstar$ CID [77] R-50-FPN 29.2 28.3
$\bigstar$ CID [77] HRNet-w32 44.9 44.0